OpenStudy (anonymous):

the age mrs Flores is 8 yrs more than three times her daughter's age. Find the age of each if ten years from now mrs Flores' age will be 5 yrs. more than twice her daughter's age?

OpenStudy (anonymous):

set equations

OpenStudy (anonymous):

Mrs. Flores daughter's age can be thought of as an unknown, x. Mrs. Flores age can be thought of as y. 3x + 8 = y (three times the daughter's age plus 8 is equal to Mrs. Flores age) In ten years, the situation changes to: y + 10 = 2(x+10) + 5 y + 10 is Mrs. Flore's age in 10 years x + 10 is her daughter's age in 10 years in ten years, 2 times her daughter's age plus five will equal her age. This equation simplifies to y = 2x + 15 Now, you have a system of equations 3x + 8 = y and y = 2x + 15 If both equations are equal to y, then 3x + 8 = 2x + 15 You can simplify to find out the value of x. Then substitute in to find the value of y.

OpenStudy (anonymous):

x= 7???

OpenStudy (anonymous):

Yes, that is what I got. Which makes Mrs. Flores 29 to begin with. 10 years later, they are 17 and 39.

OpenStudy (anonymous):

better soln: 3x+8+10= 2(x+10)+5